OpenStudy (anonymous):

How do I set this up so that I get: 7/2 and -6/5? 10x^2-23x-42=0

OpenStudy (shamim):

i can try to solve it

OpenStudy (shamim):

\[10x ^{2}-23x-42=0\]

OpenStudy (shamim):

or,\[10x ^{2}-35x+12x-42=0\]

OpenStudy (shamim):

or,\[5x(2x-7)+6(2x-7)=0\]

OpenStudy (shamim):

or,\[(2x-7)(5x+6)=0\]

OpenStudy (shamim):

anyway r u following me

OpenStudy (anonymous):

Does it always have to be broken down?

OpenStudy (shamim):

ya almost always

OpenStudy (shamim):

i think u r enough to do the rest of the math

OpenStudy (anonymous):

How do you know which numbers to use when breaking it down, or does it not matter?

OpenStudy (shamim):

ya its matter

OpenStudy (shamim):

u need to factorise 10 and 42 first

OpenStudy (shamim):

wt u will get after factorise

OpenStudy (shamim):

u will get 5,2,2,3,7

OpenStudy (shamim):

i got 35 frm the multiplication of 5 and 7

OpenStudy (shamim):

and i got 12 frm the multiplication of 2,2,3

OpenStudy (shamim):

then i hv to subtract 12 frm 35 for getting -23

OpenStudy (shamim):

-35x+12x=-23x

OpenStudy (shamim):

and -35x*12x=(10x^2)(-42)

OpenStudy (anonymous):

I get it now. I just have to practice more.